# Nested subspace learning with flags

### v1

@@ 4.4.2 Nestedness experiments for spectral clustering
We consider a 3D extension of the classical two-moons dataset for clustering with $n=100$. The ambient dimension is $p=3$ and the intrinsic dimensions that we try are $q_{1:2}=(1,2)$. We adopt a pre-processing strategy similar to Lu et al. (2016); Wang et al. (2017), i.e. we compute the affinity matrix $W\in\mathbb{R}^{n\times n}$ using an exponential kernel with standard deviation being the median of the pairwise Euclidean distances between samples. Then we compute the normalized Laplacian matrix $L=I_{n}-D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ where $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix with diagonal elements $D_{ii}=\sum_{j=1}^{n}w_{ij}$. We run Algorithm 1 on Grassmann manifolds to solve the sparse optimization problem (21), successively for $q_{1}=1$ and $q_{2}=2$. Then we plot the projections of the data points onto the optimal subspaces. We compare them to the nested projections onto the optimal flag output by running Algorithm 1 on $\operatorname{Fl}(3,(1,2))$ to solve (22). The results are shown in Figure 6.
Figure 6: Illustration of the nestedness issue in sparse spectral clustering. Given a dataset with two half-moon-shaped clusters, we plot its projection onto the optimal 1D subspace and 2D subspace obtained by solving the associated Grassmannian optimization problem (21) or flag optimization problem (22). We can see that the Grassmann representations are not nested, while the flag representations are nested and much better clustered. The last row corresponds to dividing the 2D embeddings by their norms, as commonly done in spectral clustering (Ng et al., 2001).
Figure 7: Illustration of the nestedness issue in sparse spectral clustering on the breast cancer dataset. For $q_{k}\in q_{1:d}:=(1,2,\dots,99)$, we solve the Grassmannian optimization problem (21) on $\operatorname{Gr}(100,q_{k})$ and plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ (left) and explained variances ${\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}}XX^{\top})}/{\operatorname{tr}(XX^{\top})}$ (right) as a function of $k$. We compare those quantities to the ones obtained by solving the flag optimization problem (22). We can see that the Grassmann-based method is highly non-nested while the flag of subspaces seems to be coherent with the cluster structure: the first axis is for the orange cluster while the second axis is for the blue cluster.
@@
Second, we consider the breast cancer dataset (Wolberg et al., 1995). It contains $569$ samples—from which we extract a subset of $100$ samples for computational time—with $30$ numerical attributes and two classes. One has $n=100$, $p=30$ and $C=2$. Then we run the steepest descent algorithm 1 to solve the SSC problem (22). We choose the full signature $q_{1:99}=(1,2,\dots,99)$ and compare the output flag to the individual subspaces output by running optimization on $\operatorname{Gr}(p,q_{k})$ for $q_{k}\in q_{1:d}$. We perform scatter plots in $1D$ and $2D$ and subspace error computations as a function of the $q_{k}$. The results are illustrated in Figure 7. The subspace angle plot tells us that the Grassmann spectral clustering yields highly non-nested subspaces, while the flags are by nature nested. The scatter plots show that the Grassmann representations are totally inconsistent, while the flag representations are consistent with the cluster structures.
Third, we perform some classification experiments on the optimal subspaces for different datasets. For different datasets, we run the optimization problems on $\operatorname{Fl}(p,q_{1:d})$, then project the data onto the different subspaces in $\mathcal{S}_{1:d}^{*}$, normalize the rows of their Stiefel representative $U_{k}^{*}\in\operatorname{St}(p,q_{k})$ to have unit Euclidean length and run a k-means algorithm with the number of clusters being equal to the number of classes. We report the clustering results in Table 2.
Table 2: Clustering results for the SSC problem on real datasets. For each method (Gr: Grassmann optimization (21), Fl: flag optimization (22)), we give the Rand index metric (Rand, 1971) of the projected-predictions with respect to the true cluster labels.
dataset $n$ $p$ $q_{1:d}$ Gr Fl
iris $99$ $4$ $(1,3,5)$ $0.8846$ $0.9371$
breast $100$ $30$ $(1,2,5)$ $0.5533$ $0.8181$
wine $99$ $13$ $(1,3,5)$ $0.6910$ $0.7108$
digits $100$ $64$ $(1,2,5,10)$ $0.9743$ $0.9820$
We can see that the flag approach yields more discriminating subspaces than the Grassmann approach. This can be explained by the multilevel nature of flags.
@@ 4.4.3 Discussion on SSC optimization
Similarly as in the previous examples, we can see that the steepest descent algorithm 1 might not be the most suited to solve the optimization problem (22), notably due to the $\ell_{1}$ penalization. Contrary to the RSR case in which we want to avoid the non-differentiable points (cf. subsubsection 4.2.3), here we would ideally like to attain those points, since they encode the desired sparsity of the Laplacian matrix. Therefore, we believe that more standard methods taking into account the $\ell_{1}$ norm to enforce the hierarchical sparsity of the flag of subspaces would be more adapted.
@@ 4.5.1 Domain adaptation
In machine learning, it is often assumed that the training and test datasets follow the same distribution. However, some domain shift issues—where training and test distributions are different—might arise, notably if the test data has been acquired from a different source (for instance a professional camera and a phone camera) or if the training data has been acquired a long time ago. Domain adaptation is an area of machine learning that deals with domain shifts, usually by matching the training and test distributions—often referred to as source and target distributions—before fitting a classical model (Farahani et al., 2021). A large body of works (called “subspace-based”) learn some intermediary subspaces between the source and target data, and perform the inference for the projected target data on these subspaces. The sampling geodesic flow (Gopalan et al., 2011) first performs a geodesic interpolation on Grassmannians between the source and target subspaces, then projects both datasets on (a discrete subset of) the interpolated subspaces, which results in a new representation of the data distributions, that can then be given as an input to a machine learning model. The higher the number of intermediary subspaces, the better the approximation, but the larger the dimension of the representation. The celebrated geodesic flow kernel (Boqing Gong et al., 2012) circumvents this issue by integrating the projected data onto the continuum of interpolated subspaces. This yields an inner product between infinite-dimensional embeddings that can be computed explicitly and incorporated in a kernel method for learning. The domain invariant projection (Baktashmotlagh et al., 2013) learns a domain-invariant subspace that minimizes the maximum mean discrepancy (MMD) (Gretton et al., 2012) between the projected source $X_{s}:=[x_{s1}|\dots|x_{sn_{s}}]\in\mathbb{R}^{p\times n_{s}}$ and target distributions $X_{t}:=[x_{t1}|\dots|x_{tn_{t}}]\in\mathbb{R}^{p\times n_{t}}$:
$\underset{U\in\operatorname{St}(p,q)}{\operatorname{argmin}}\,\operatorname{MMD}^{2}(U^{\top}X_{s},U^{\top}X_{t}),$ (24)
where
$\operatorname{MMD}(X,Y)=\left\lVert\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})-\frac{1}{m}\sum_{i=1}^{m}\phi(y_{i})\right\rVert_{\mathcal{H}}.$ (25)
This can be rewritten, using the Gaussian kernel function $\phi(x)\colon y\mapsto\exp\left(-\frac{x^{\top}y}{2\sigma^{2}}\right)$, as
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}\exp\left(-\frac{(x_{si}-x_{sj})^{\top}\Pi_{\mathcal{S}}(x_{si}-x_{sj})}{2\sigma^{2}}\right)\\ +\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}\exp\left(-\frac{(x_{ti}-x_{tj})^{\top}\Pi_{\mathcal{S}}(x_{ti}-x_{tj})}{2\sigma^{2}}\right)\\ -\frac{2}{n_{s}n_{t}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}\exp\left(-\frac{(x_{si}-x_{tj})^{\top}\Pi_{\mathcal{S}}(x_{si}-x_{tj})}{2\sigma^{2}}\right).$ (26)
The flag trick applied to the domain invariant projection problem (26) yields:
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}\exp\left(-\frac{(x_{si}-x_{sj})^{\top}\Pi_{\mathcal{S}_{1:d}}(x_{si}-x_{sj})}{2\sigma^{2}}\right)\\ +\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}\exp\left(-\frac{(x_{ti}-x_{tj})^{\top}\Pi_{\mathcal{S}_{1:d}}(x_{ti}-x_{tj})}{2\sigma^{2}}\right)\\ -\frac{2}{n_{s}n_{t}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}\exp\left(-\frac{(x_{si}-x_{tj})^{\top}\Pi_{\mathcal{S}_{1:d}}(x_{si}-x_{tj})}{2\sigma^{2}}\right),$ (27)
and can be rewritten as:
$\underset{U_{1:d}\in\operatorname{St}(p,q)}{\operatorname{argmin}}\,\frac{1}{{n_{s}}^{2}}\sum_{i,j=1}^{n_{s}}\exp\left(-\sum_{k=1}^{d}\frac{d+1-k}{d}\frac{\left\lVert{U_{k}}^{\top}(x_{si}-x_{sj})\right\rVert_{2}^{2}}{2\sigma^{2}}\right)\\ +\frac{1}{{n_{t}}^{2}}\sum_{i,j=1}^{n_{t}}\exp\left(-\sum_{k=1}^{d}\frac{d+1-k}{d}\frac{\left\lVert{U_{k}}^{\top}(x_{ti}-x_{tj})\right\rVert_{2}^{2}}{2\sigma^{2}}\right)\\ -\frac{2}{{n_{s}}{n_{t}}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}\exp\left(-\sum_{k=1}^{d}\frac{d+1-k}{d}\frac{\left\lVert{U_{k}}^{\top}(x_{si}-x_{tj})\right\rVert_{2}^{2}}{2\sigma^{2}}\right).$ (28)
@@
Some experiments similar to the ones of Baktashmotlagh et al. (2013) can be performed. For instance, one can consider the benchmark visual object recognition dataset of Saenko et al. (2010), learn nested domain invariant projections, fit some support vector machines to the projected source samples at increasing dimensions, and then perform soft-voting ensembling by learning the optimal weights on the target data according to Equation (19).
@@ 4.5.2 Low-rank decomposition
Many machine learning methods involve finding low-rank representations of a data matrix.
This is the case of matrix completion (Candès and Recht, 2012) problems where one looks for a low-rank representation of an incomplete data matrix by minimizing the discrepancy with the observed entries, and which finds many applications including the well-known Netflix problem. Although its most-known formulation is as a convex relaxation, it can also be formulated as an optimization problem on Grassmann manifolds (Keshavan et al., 2010; Boumal and Absil, 2011) to avoid optimizing the nuclear norm in the full space which can be of high dimension. The intuition is that a low-dimensional point can be described by the subspace it belongs to and its coordinates within this subspace. More precisely, the SVD-based low-rank factorization $M=UW$, with $M\in\mathbb{R}^{p\times n}$, $U\in\operatorname{St}(p,q)$ and $W\in\mathbb{R}^{q\times n}$ is orthogonally-invariant—in the sense that for any $R\in\mathcal{O}(q)$, one has $(UR)(R^{\top}W)=UW$. One could therefore apply the flag trick to such problems, with the intuition that we would try low-rank matrix decompositions at different dimensions. The application of the flag trick would however not be as straightforward as in the previous problems since the subspace-projection matrices $\Pi_{\mathcal{S}}:=UU^{\top}$ do not appear explicitly, and since the coefficient matrix $W$ also depends on the dimension $q$.
Many other low-rank problems can be formulated as a Grassmannian optimization. Robust PCA (Candès et al., 2011) looks for a low rank + sparse corruption factorization of a data matrix. Subspace Tracking (Balzano et al., 2010) incrementally updates a subspace from streaming and highly-incomplete observations via small steps on Grassmann manifolds.
4.5.3 Linear dimensionality reduction
Finally, many other general dimension reduction algorithms—referred to as linear dimensionality reduction methods (Cunningham and Ghahramani, 2015)—involve optimization on Grassmannians. For instance, linear dimensionality reduction encompasses the already-discussed PCA and LDA, but also many other problems like multi-dimensional scaling (Torgerson, 1952), slow feature analysis (Wiskott and Sejnowski, 2002), locality preserving projections (He and Niyogi, 2003) and factor analysis (Spearman, 1904).
@@ 5 Discussion
We introduced a general method to make subspace learning problems multilevel and nested. The nestedness issue of Grassmannian-based methods was clearly demonstrated on a variety of machine learning problems, yielding inconsistent scatter plots and paradoxical learning curves. The flag trick was shown to both solve the nestedness issue and provide solutions that are close from the original objective of the problems. As a positive side-effect, the flag trick sometimes learned substantially better subspaces than the original subspace problems, due to its multilevel nature which enables to aggregate information from several dimensions. When combined with ensemble learning, the flag trick showed improvement over subspace-based predictors that work at a fixed dimension, and raised interesting dimension-blending perspectives that challenge the manifold hypothesis. The code for reproducing the experiments can be found at https://github.com/tomszwagier/flag-trick.
@@
One major limitation of this paper is the lack of in-depth study of the specific interest of the flag trick to each subspace learning method. We indeed focus on general practical results—the nestedness of flag-based methods with respect to classical subspace methods and the interest of ensembling dimensions instead of considering each dimension individually—and not on some application-specific questions, like the robustness of the flag to outliers for robust subspace recovery, the quality of the embedding for trace ratio problems, and the quality of the clustering for spectral clustering problems. We therefore strongly encourage the specialists to investigate with us the theoretical and practical interest of the flag trick for their subspace problems. We also encourage the Riemannian optimization community to develop with us new optimization algorithms on flag manifolds that are adapted to each subspace learning problem—for instance with robustness or sparsity-inducing properties. To that extent, we end this paper with a list of perspectives.
First, one should perform in-depth experiments to show the interest and the limitations of the flag trick for each specific subspace learning problem among—but not limited to—the list of examples developed in section 4, as well as their local, kernel and tangent space generalizations. Second, one should experiment with more complex and more efficient optimization algorithms on flag manifolds such as the ones developed in Appendix (IRLS in Appendix A and Newton in Appendix B) or the ones proposed in Ye et al. (2022); Zhu and Shen (2024) and develop new ones, like ADMM for sparse optimization. Third, we derived a very general principle for transitioning from fixed-dimension subspace methods to multi-dimensional flag methods, but this principle could be revisited specifically for each problem. This includes the problems that are not specifically formulated as a Grassmannian optimization, as long as they somewhat involve low-dimensional subspaces, like in domain adaptation (Gopalan et al., 2011; Boqing Gong et al., 2012) and sparse subspace clustering (Elhamifar and Vidal, 2013). Fourth, the question of how to choose the flag signature (i.e. the sequence of dimensions to try) remains open. A first step for PCA has been achieved with the principal subspace analysis (Szwagier and Pennec, 2024), where the dimensions are chosen based on the relative eigengaps of the covariance matrix. Other generative models should be designed specifically for each subspace learning problem in order to develop some appropriate guidelines to select the flag signature.
@@
This work was supported by the ERC grant #786854 G-Statistics from the European Research Council under the European Union’s Horizon 2020 research and innovation program and by the French government through the 3IA Côte d’Azur Investments ANR-19-P3IA-0002 managed by the National Research Agency.

### v2

@@ 4.4.2 Nestedness Experiments for Spectral Clustering
We run some nestedness experiments for the sparse spectral clustering problem. First, we consider a 3D extension of the classical two-moons dataset for clustering with $n=100$. The ambient dimension is $p=3$ and the intrinsic dimensions that we try are $q_{1:2}=(1,2)$. We adopt a pre-processing strategy similar to lu_convex_2016; wang_grassmannian_2017: we compute the affinity matrix $W\in\mathbb{R}^{n\times n}$ using an exponential kernel with standard deviation being the median of the pairwise Euclidean distances between samples. Then we compute the normalized Laplacian matrix $L=I_{n}-D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ where $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix with diagonal elements $D_{ii}=\sum_{j=1}^{n}w_{ij}$. We run Algorithm 1 on Grassmann manifolds to solve the sparse optimization problem (22), successively for $q_{1}=1$ and $q_{2}=2$. Then we plot the projections of the data points onto the optimal subspaces. We compare them to the nested projections onto the optimal flag output by running Algorithm 1 on $\operatorname{Fl}(3,(1,2))$ to solve (23). The results are shown in Figure 7.
Figure 7: Illustration of the nestedness issue in sparse spectral clustering. Given a dataset with two half-moon-shaped clusters, we plot its projection onto the optimal 1D subspace and 2D subspace obtained by solving the associated Grassmannian optimization problem (22) or flag optimization problem (23). We can see that the Grassmann representations are not nested, while the flag representations are nested and much better clustered. The last row corresponds to dividing the 2D embeddings by their norms, as commonly done in spectral clustering (ng_spectral_2001).
Figure 8: Illustration of the nestedness issue in sparse spectral clustering on the breast cancer dataset. On the left, for $q_{k}\in(1,2,\dots,99)$, we solve the Grassmannian optimization problem (22) on $\operatorname{Gr}(100,q_{k})$ and plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ as a function of $k$. On the right, we plot the projection of the dataset onto the output 1D and 2D subspaces. We compare those quantities to the ones obtained by solving the flag optimization problem (23). We can see that the Grassmann-based method is highly non-nested while the flag-tricked one is not only nested but also seems to be coherent with the cluster structure: the first axis is for the red cluster while the second axis is for the blue cluster.
@@
Second, we consider the breast cancer dataset (wolberg_breast_1995). It contains $569$ samples—from which we extract a subset of $100$ samples for computational time—with $30$ numerical attributes and two classes. One has $n=100$, $p=30$ and $C=2$. Then we run the steepest descent algorithm 1 to solve the SSC problem (23). We choose the full signature $q_{1:99}=(1,2,\dots,99)$ and compare the output flag to the individual subspaces output by running optimization on $\operatorname{Gr}(n,q_{k})$ for $q_{k}\in q_{1:d}$. We perform scatter plots in 1D and 2D and subspace error computations as a function of the $q_{k}$. The results are illustrated in Figure 8. The subspace angle plot tells us that the Grassmann spectral clustering yields highly non-nested subspaces, while the flags are by nature nested. The scatter plots show that the Grassmann representations are totally inconsistent, while the flag representations are consistent with the cluster structures.
@@ 4.4.3 Discussion on SSC Optimization
Similarly to the previous examples, we can see that the steepest descent algorithm 1 might not be the most suited to solve the optimization problem (23), notably due to the $\ell^{1}$-penalization. Contrary to robust subspace recovery, where we want to avoid the non-differentiable points (cf. Section 4.2.3), here we would ideally like to exactly attain those points, since they encode the desired sparsity of the Laplacian matrix. Therefore, we believe that sparsity-inducing optimization algorithms (bach_optimization_2012) on flag manifolds would be more adapted than the steepest descent we use.
@@ 4.5.1 Domain Adaptation
In machine learning, it is often assumed that the training and test datasets follow the same distribution. However, some domain shift issues—where training and test distributions are different—might arise, notably if the test data has been acquired from a different source (for instance a professional camera and a phone camera) or if the training data has been acquired a long time ago. Domain adaptation is an area of machine learning that deals with domain shifts, usually by matching the training and test distributions—often referred to as source and target distributions—before fitting a classical model (farahani_brief_2021). A large body of works (called “subspace-based”) learn some intermediary subspaces between the source and target data, and perform the inference for the projected target data on these subspaces. The sampling geodesic flow (gopalan_domain_2011) first performs a geodesic interpolation on Grassmannians between the source and target subspaces, then projects both datasets on (a discrete subset of) the interpolated subspaces, which results in a new representation of the data distributions, that can then be given as an input to a machine learning model. The higher the number of intermediary subspaces, the better the approximation, but the larger the dimension of the representation. The celebrated geodesic flow kernel (boqing_gong_geodesic_2012) circumvents this issue by integrating the projected data onto the continuum of interpolated subspaces. This yields an inner product between infinite-dimensional embeddings that can be computed explicitly and incorporated in a kernel method for learning. The domain invariant projection (baktashmotlagh_unsupervised_2013) learns a domain-invariant subspace that minimizes the maximum mean discrepancy (MMD) (gretton_kernel_2012) between the projected source $X_{s}\coloneqq\begin{bmatrix}x_{s1}&\cdots&x_{sn_{s}}\end{bmatrix}\in\mathbb{R}^{p\times n_{s}}$ and target distributions $X_{t}\coloneqq\begin{bmatrix}x_{t1}&\cdots&x_{tn_{t}}\end{bmatrix}\in\mathbb{R}^{p\times n_{t}}$: $\operatorname{argmin}_{U\in\operatorname{St}(p,q)}\operatorname{MMD}^{2}(U^{\top}X_{s},U^{\top}X_{t})$, where $\operatorname{MMD}(X,Y)=\left\lVert\frac{1}{n}\sum_{i=1}^{n}\phi(x_{i})-\frac{1}{m}\sum_{i=1}^{m}\phi(y_{i})\right\rVert_{\mathcal{H}}$. Following baktashmotlagh_unsupervised_2013, we can rewrite the unsupervised domain adaptation problem as:
$\operatorname*{arg\,min}_{\mathcal{S}\in\operatorname{Gr}(p,q)}\,\left(\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}\exp\left(-\frac{{z^{s}_{ij}}^{\top}\Pi_{\mathcal{S}}z^{s}_{ij}}{2\sigma^{2}}\right)+\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}\exp\left(-\frac{{z^{t}_{ij}}^{\top}\Pi_{\mathcal{S}}z^{t}_{ij}}{2\sigma^{2}}\right)\right.\\ \left.-\frac{2}{n_{s}n_{t}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}\exp\left(-\frac{{z^{st}_{ij}}^{\top}\Pi_{\mathcal{S}}z^{st}_{ij}}{2\sigma^{2}}\right)\right),$ (25)
where $z^{s}_{ij}\coloneqq x_{si}-x_{sj}$, $z^{t}_{ij}\coloneqq x_{ti}-x_{tj}$ and $z^{st}_{ij}\coloneqq x_{si}-x_{tj}$. One can easily apply the flag trick (11) to this problem to make domain adaptation multilevel:
$\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\,\left(\frac{1}{n_{s}^{2}}\sum_{i,j=1}^{n_{s}}\exp\left(-\frac{{z^{s}_{ij}}^{\top}\Pi_{\mathcal{S}_{1:d}}z^{s}_{ij}}{2\sigma^{2}}\right)+\frac{1}{n_{t}^{2}}\sum_{i,j=1}^{n_{t}}\exp\left(-\frac{{z^{t}_{ij}}^{\top}\Pi_{\mathcal{S}_{1:d}}z^{t}_{ij}}{2\sigma^{2}}\right)\right.\\ \left.-\frac{2}{n_{s}n_{t}}\sum_{i=1}^{n_{s}}\sum_{j=1}^{n_{t}}\exp\left(-\frac{{z^{st}_{ij}}^{\top}\Pi_{\mathcal{S}_{1:d}}z^{st}_{ij}}{2\sigma^{2}}\right)\right).$ (26)
@@
Some experiments similar to the ones of baktashmotlagh_unsupervised_2013 can be performed. For instance, one can consider the benchmark visual object recognition dataset of saenko_adapting_2010, learn nested domain invariant projections, fit some support vector machines to the projected source samples at increasing dimensions, and then perform soft-voting ensembling by learning the optimal weights on the target data according to Equation (20).
@@ 4.5.2 Low-Rank Decomposition
Many machine learning methods involve finding low-rank representations of a data matrix. This is the case of matrix completion (candes_exact_2012) where one looks for a low-rank representation of an incomplete data matrix by minimizing the discrepancy with the observed entries, and which finds many applications including the well-known Netflix problem. Although its most-known formulation is as a convex relaxation, it can also be formulated as an optimization problem on Grassmann manifolds (keshavan_matrix_2010; boumal_rtrmc_2011) to avoid optimizing the nuclear norm in the full space which can be of high dimension. The intuition is that a low-dimensional point can be described by the subspace it belongs to and its coordinates within this subspace. More precisely, the SVD-based low-rank factorization $M=UW$, with $M\in\mathbb{R}^{p\times n}$, $U\in\operatorname{St}(p,q)$ and $W\in\mathbb{R}^{q\times n}$ is orthogonally-invariant—in the sense that for any $R\in\mathcal{O}(q)$, one has $(UR)(R^{\top}W)=UW$. One could therefore apply the flag trick to such problems, with the intuition that we would try low-rank matrix decompositions at different dimensions. The application of the flag trick would however not be as straightforward as in the previous problems since the subspace-projection matrices $\Pi_{\mathcal{S}}\coloneqq UU^{\top}$ do not appear explicitly, and since the coefficient matrix $W$ also depends on the dimension $q$.
Many other low-rank problems can be formulated as a Grassmannian optimization. Robust PCA (candes_robust_2011) looks for a low rank + sparse corruption factorization of a data matrix. Subspace Tracking (balzano_online_2010) incrementally updates a subspace from streaming and highly-incomplete observations via small steps on Grassmann manifolds.
4.5.3 Other subspace learning problems
Finally, many other general machine learning problems involve optimization on Grassmannians. For instance, linear dimensionality reduction (cunningham_linear_2015) encompasses the already-discussed PCA and LDA, but also many other problems like multi-dimensional scaling (torgerson_multidimensional_1952), slow feature analysis (wiskott_slow_2002) and locality preserving projections (he_locality_2003). All these machine learning problems involving optimization on Grassmannians can, likewise, be turned into multilevel problems via the flag trick.
@@ 5 Discussion
We introduced a general principle to make subspace learning problems multilevel and nested. The non-nestedness issue of Grassmannian-based methods was clearly demonstrated on a variety of machine learning problems, yielding inconsistent scatter plots and paradoxical learning curves. The flag trick was shown to both solve this nestedness issue and output filtrations of subspaces that meet the original goals of the problems. As a positive side-effect, the flag trick sometimes even output substantially-better subspaces than the original subspace problems, due to its multilevel nature which enables to aggregate information from several dimensions. When combined with ensemble learning, the flag trick showed an overall improvement over the subspace-based predictors that work at a fixed dimension, and it raised some interesting dimension-blending perspectives that question the manifold hypothesis. The code for the experiments can be found at https://github.com/tomszwagier/flag-trick.
@@
One major limitation of this paper is the lack of in-depth study of the specific interest of the flag trick to each subspace learning method. We indeed focus on general practical results—the nestedness of flag-based methods with respect to classical subspace methods and the interest of ensembling dimensions instead of considering each dimension individually—and not on some application-specific questions—like the robustness of the flag to outliers for robust subspace recovery, the quality of the embedding for trace ratio problems, and the quality of the clustering for spectral clustering problems. Therefore, this paper should rather be seen as an invitation for the different communities to explore the multilevel extensions of subspace learning, than as the unique solution to a long-standing issue in machine learning. To that extent, we end this paper with a list of perspectives.
First, one should perform in-depth experiments to show the interest and the limitations of the flag trick for each specific subspace learning problem among—but not limited to—the list of examples developed in Section 4. Throughout this paper, we focused on linear flags, but nonlinear extensions could be considered to improve expressivity. Beyond the kernel trick considered in Section B.2 for nonlinear graph embedding, a general methodology based on data mappings to linear spaces—such as the latent space of a deep generative model or the tangent space at the Fréchet mean of a manifold-valued dataset—could be simply implemented. More advanced methodologies relying on nested submanifolds, borrowed from the geometric statistics community (see Section 1), could be investigated as optimization problems on flag manifolds; typical examples are the principal nested spheres from jung_analysis_2012, which can be seen as intersections between spheres and (linear) hyperplanes.
Second, one should experiment with more complex and more efficient optimization algorithms on flag manifolds such as the ones developed in Appendix (IRLS in Appendix A and Newton in Appendix B) or the ones proposed in ye_optimization_2022, nguyen_closed-form_2022 and zhu_practical_2024 and develop new ones that are specifically adapted to the properties of the objective function. We invite the interested researchers from the manifold optimization community to test their existing algorithms (and develop new ones) on our statistically-grounded optimization problems (13), (18), (23), (26) on flag manifolds.
Third, we derived a very general principle for transitioning from fixed-dimension subspace methods to increasing-dimension flag methods, but this principle could be revisited specifically for each problem. This includes the problems that are not specifically formulated as a Grassmannian optimization, as long as they somewhat involve low-dimensional subspaces, like in domain adaptation (gopalan_domain_2011; boqing_gong_geodesic_2012) and sparse subspace clustering (elhamifar_sparse_2013).
Fourth, we leave the question of the choice of flag signature (i.e. the sequence of dimensions to try) open in this paper (see, notably, Remark 12). A first step for PCA in the Gaussian case has been achieved with the principal subspace analysis of szwagier_curse_2025, where the dimensions are chosen based on the relative eigengaps of the covariance matrix. Other generative models could be designed specifically for each subspace learning problem, and the recent penalty of szwagier_eigengap_2025 on the spectral gaps could be added to automatically select the flag signature.
@@
The authors would like to thank Pierre-Antoine Absil for his insightful comments on the flag trick and its applications. This work was supported by the ERC grant #786854 G-Statistics from the European Research Council under the European Union’s Horizon 2020 research and innovation program and by the French government through the 3IA Côte d’Azur Investments ANR-23-IACL-0001 managed by the National Research Agency.